OpenStudy (anonymous):

7*7^2/7^9 use properties to simplify

OpenStudy (anonymous):

7^3/7^9. now use the propeties...

OpenStudy (diyadiya):

\[\frac{7 \times 7^2}{7^9}\] \[\huge x^m \times x^n= x^{m+n}\]

OpenStudy (anonymous):

what do you do to the exponents when you have quotient o powers?

OpenStudy (diyadiya):

\[\huge \frac{x^m}{x^n}= x^{m-n}\]

OpenStudy (anonymous):

what would be the answer

OpenStudy (diyadiya):

by using this property \[\huge x^m \times x^n= x^{m+n}\] What will be \[7 \times 7^2 =?\]

OpenStudy (diyadiya):

\[\text{Remember 7 is } 7^1\]

OpenStudy (anonymous):

@Diyadiya how do you write that kind of multiply with latex ?

OpenStudy (diyadiya):

write multiply as times For example 3 times 7

OpenStudy (anonymous):

\[\LARGE \frac{4}{0\times 7}= 5 \quad \quad O _ O^OO\] Now I get it :F

OpenStudy (diyadiya):

@9013955 ??